# Podolsky quantum electrodynamics for strongly coupled Dirac fermions in (2+1)D

### v1

@@ Abstract
We investigate Podolsky quantum electrodynamics (POQED), a higher-derivative extension of QED in (3+1)D, and perform its dimensional reduction to (2+1)D by confining the Dirac current to a plane while allowing the gauge field to propagate out of the plane. The resulting model, which we call Pseudo Generalized QED (PGQED), is minimally coupled to massless Dirac fermions. In the strong-coupling regime, we show that a dynamical mass is generated through approximate solutions of the Schwinger-Dyson equation, leading to chiral symmetry breaking and modifications of the fermion dispersion relation. We derive an analytic critical coupling $\alpha_{c}(\mu)$ and flavors critical number $N_{c}(\mu)$, dependent on the Podolsky parameter $\mu$ and the ultraviolet cutoff $\Lambda$. These analytical results are found to be consistent with the numerical analysis. Finally, we discuss connections to graphene, estimating a range for $\mu$ in the ultrarelativistic limit.
@@ I INTRODUCTION
The GQED, proposed by Boris Podolsky in 1942, is a model that introduces a higher-order derivative term into the standard quantum electrodynamics (QED), yielding a characteristic parameter $a$ with units of inverse of mass, called the Podolsky parameter [15]. Whenever this parameter vanishes, one recovers Maxwell’s theory. This GQED preserves both Lorentz and gauge symmetries while describing a massive mode for the gauge field, beyond the standard massless photon [16]. Furthermore, GQED provides a finite potential for a point-like particle at short distances, unlike electrons within conventional QED [15]. Hence, one expects a better UV behavior of the model in comparison to QED, which may have consequences for renormalizability of the theory. The dimensional reduction of GQED has not been considered until now. In Ref. [17] a similar model has been considered within classical field theory.
@@
The dynamical mass generation implies a chiral symmetry breaking in (2+1)D whenever we consider Dirac fermions in the four-rank representation of Dirac matrices [1, 3]. Hence, for strong interactions, we consider a suitable set of approximations to solve the Schwinger-Dyson equations. In particular, we find nontrivial solutions for the mass function $\Sigma(p)$ of the full electron propagator, using both the rainbow-quenched and rainbow-unquenced approximations. In general, when $\Sigma(p)\neq 0$, we have a dynamically generated mass. For $\Sigma(p\rightarrow 0)=M$, which is a quite simplistic solution, it follows that $M$ is the pole of the full fermion propagator, therefore, it would be the dynamical mass of the electron. Furthermore, we show that $M$ corresponds to a small fraction of the UV cutoff $\Lambda$, which plays an important role in our model. This parameter can be related to the lattice constant $a_{L}\approx 10^{-10}$m of graphene, hence $\Lambda\approx 1$eV which works as an upper energy limit for our Dirac cones [18]. Indeed, for energies above this value one would observe corrections to the Dirac cones and, therefore, the Dirac approximation for electrons in graphene becomes less reliable [5]. Here, we consider $\Lambda$ to be a large energy value. We also consider a full Lorentz invariant matter field, which simplifies the calculation of the electron self-energy. This can also be realized in graphene in the low-density limit of charge carriers, as discussed in Ref. [18]. Improved approximations show that there exist a critical coupling constant $\alpha_{c}$ and a critical number of fermions $N_{c}$ that separate the symmetry and broken phases. These parameters bound the values of the fine-structure constant $\alpha\propto e^{2}/4\pi$ ($\alpha>\alpha_{c}$) and the number of degrees of freedom NN ($N<N_{c}$). Finally, we also describe the relevance of the Podolsky parameter $\mu=a^{-1}$ for the onset of this quantum phase transition.
The outline of this paper is following: In Sec. II, we derive our model and its Feynman rules. In Sec. III, we discuss the Schwinger-Dyson equation for our model. In Sec. IV and V, we calculate the mass function using a suitable set of approximations. In Sec. VI, we discuss our main results and their consequences. We also include one Appendix A where we show details about the dynamical mass generation and Appendix B where we show the numerical results for the mass function and the renormalization of the wavefunction of the electron.
@@ II THE MODEL
For small distances $r\ll a$, the potential is finite and given by $e^{2}/4\pi a$ at lowest-order approximation, which is a feature of GQED. For larger distances $r\gg a$, we obtain the Coulomb potential as in QED in (3+1)D. On the other hand, for intermediate values $r\approx a$, we have a Coulomb potential with a screened electric charge. In Fig. 1, we plot Eq. (4) using an artificial value for $a$. These results mostly concern the classical physics of GQED. From now on, we shall consider quantum effects of the dimensional reduction of GQED.
@@ II.1 The Feynman Rules
In this section, we derive the Feyman rules for PGQED in Eq. (II) at tree level. This is mandatory to apply the so-called Schwinger-Dyson equations for calculating the full Green functions of the model, which we shall discuss later. After a standard algebra, we obtain the bare gauge-field propagator, given by
@@
where use the Landau gauge, i.e, $\xi\to+\infty$. After integrating out the gauge field in Eq. (II), it is straightforward to conclude that static charges interact through the potential $V(r)$ in Eq. (4), which is obtained from the Fourier transform of $\Delta^{0}_{00}(p_{0}=0,\textbf{p})$ [20]. Note that whether $\mu\rightarrow\infty$ (and equivalently $a\rightarrow 0$), we recover the PQED model as expected. Furthermore, in contrast to standard GQED, we can not conclude that this propagator in Eq. (5) describes either a massive or massless particle. Indeed, due to the square root in the gauge-field propagator, it does not have a pole and it needs a branch cut in order to become a single-valued function in the complex plane of $p$. Some important features of this propagator have been discussed in Ref. [17] at classical level.
@@ IV RAINBOW-QUENCHED APPROXIMATION
which is a differential equation for the mass function $\Sigma(p)$. This must also satisfy both the IR and UV conditions, given by Eqs. (29) and (30), respectively. Even after our set of approximations, Eq. (31) still is a nonlinear differential equation due to the denominator $p^{2}+\Sigma^{2}(p)$. This is a typical property of the Schwinger-Dyson equation, where the full propagator is written as a function of itself. Next, we propose a few approximations in order to circumvent this problem.
@@ IV.1 Approximations: $p\to 0$
$\frac{\pi}{2\alpha}$ $=\frac{1}{2}\ln{\frac{M^{2}+\Lambda^{2}}{M^{2}}}-\ln{\frac{\sqrt{\mu^{2}+\Lambda^{2}}+\Lambda}{\mu}}+$
$+\frac{|M|}{\sqrt{\mu^{2}-M^{2}}}\arctan{\frac{\Lambda\sqrt{\mu^{2}-M^{2}}}{|M|\sqrt{\mu^{2}+\Lambda^{2}}}}.$ (33)
@@
$\frac{\pi}{2\alpha}\approx\ln{\frac{\Lambda}{|M|}}-\ln{\frac{\sqrt{\mu^{2}+\Lambda^{2}}+\Lambda}{\mu}},$ (34)
@@ VI SUMMARY AND DISCUSSIONS
Recently, the author in Ref. [17] has performed a dimensional reduction in the Lee-Wick electrodynamics from (3+1)D to (2+1)D. This calculation yields the same reduced gauge-field propagator we have obtained here, given by Eq. (5) in momentum space, although the starting action in (3+1)D is slightly different from ours. However, they are dual because the electrons interact through the same static potential $V(r)\propto(1-e^{-\mu r})/r$. In Ref. [17], it is also shown that this kind of model respects the microcausality at tree level, which is a desirable feature for a well-defined field theory. Here, our main results depend on the strength of quantum correction and are likely to be the same in the context of the reduced Lee-Wick electrodynamics, given the similarity of its gauge-field propagator in comparison to the one in PGQED. Obviously, our main results obtained here only hold for an initially massless Dirac field.
@@ Acknowledgement
C. A. P. C. J. is partially supported by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior Brasil (CAPES), finance code 001. L. O. N. and V. S. A. are partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico-Brazil (CNPq-Brazil), process: 408735/2023-6 CNPq/MCTI.

### v2

@@ Abstract
We investigate generalized quantum electrodynamics (GQED), a higher-derivative extension of QED in (3+1)D. We perform its dimensional reduction to (2+1)D by confining the Dirac current to a plane while allowing the gauge field to propagate out of the plane. The resulting model, which we call Pseudo Generalized QED (PGQED), is minimally coupled to massless Dirac fermions. In the strong-coupling regime, we show that a dynamical mass is generated through approximate solutions of the Schwinger-Dyson equations, leading to chiral symmetry breaking and modifications of the fermion dispersion relation. We derive an analytic critical coupling $\alpha_{c}(\mu)$ and flavors critical number $N_{c}(\mu)$, dependent on the Podolsky parameter $\mu$ and the ultraviolet cutoff $\Lambda$. These analytical results are found to be consistent with the numerical analysis. Finally, we connect our results to graphene, estimating a range for $\mu$ in the ultrarelativistic limit and highlighting implications for two-dimensional materials.
@@ I INTRODUCTION
The GQED, proposed by Boris Podolsky in 1942, is a model that introduces a higher-order derivative term into the standard quantum electrodynamics (QED), yielding a characteristic parameter $a$ with units of inverse of mass, called the Podolsky parameter [15]. Whenever this parameter vanishes, one recovers Maxwell’s theory. This GQED preserves both Lorentz and gauge symmetries while describing a massive mode for the gauge field, beyond the standard massless photon [16]. Furthermore, GQED provides a finite potential for a point-like particle at short distances, unlike electrons within conventional QED [15]. Hence, one expects a better UV behavior of the model in comparison to QED, which may have consequences for renormalizability of the theory. Recently, some results have been discussed in literature about the GQED: (i) Gauge-fixing terms and Pauli-Villars regularization in Ref. [17]; (ii) canonical and path-integral quantization in Ref. [18], and (iii) dynamics using the light-front coordinates in Ref. [19]. The dimensional reduction of GQED has not been considered until now. Our work fills this gap by deriving PGQED and demonstrating its superior UV behavior compared to PQED. In Ref. [20], a similar model has been considered and its Higgs mechanism and unitarity at tree level have been discussed.
@@
The dynamical mass generation implies a chiral symmetry breaking in (2+1)D whenever we consider Dirac fermions in the four-rank representation of Dirac matrices [1, 3]. Hence, for strong interactions, we consider a suitable set of approximations to solve the Schwinger-Dyson equations. In particular, we find nontrivial solutions for the mass function $\Sigma(p)$ of the full electron propagator, using both the rainbow-quenched and rainbow-unquenced approximations. In general, when $\Sigma(p)\neq 0$, we have a dynamically generated mass. For $\Sigma(p\rightarrow 0)=M$, which is a quite simplistic solution, it follows that $M$ is the pole of the full fermion propagator, therefore, it would be the dynamical mass of the electron. Furthermore, we show that $M$ corresponds to a small fraction of the UV cutoff $\Lambda$, which plays an important role in our model. This parameter can be related to the lattice constant $a_{L}\approx 10^{-10}$m of graphene, hence $\Lambda\approx 1$eV which works as an upper energy limit for our Dirac cones [21]. Indeed, for energies above this value one would observe corrections to the Dirac cones and, therefore, the Dirac approximation for electrons in graphene becomes less reliable [5]. Here, we consider $\Lambda$ to be a large energy value. We also consider a full Lorentz invariant matter field, which simplifies the calculation of the electron self-energy. This can also be realized in graphene in the low-density limit of charge carriers, as discussed in Ref. [21]. In this context, our model parameters $(\Lambda,\mu,\alpha)$ can be estimated from experimental data. Improved approximations show that there exist a critical coupling constant $\alpha_{c}$ and a critical number of fermions $N_{c}$ that separate the symmetry and broken phases. These parameters bound the values of the fine-structure constant $\alpha\propto e^{2}/4\pi$ ($\alpha>\alpha_{c}$) and the number of degrees of freedom NN ($N<N_{c}$). Finally, we describe the role of the Podolsky parameter $\mu=1/a$ in triggering this quantum phase transition.
The outline of this paper is as follows: In Sec. II, we derive our model and its Feynman rules. In Sec. III, we discuss the Schwinger-Dyson equations for our model. In Sec. IV and V, we calculate the mass function using a suitable set of approximations. In Sec. VI, we discuss our main results and their consequences. We also include one Appendix A where we show details about the dynamical mass generation and Appendix B where we show the numerical results for the mass function and the renormalization of the wavefunction of the electron.
@@ II THE MODEL
For small distances $r\ll a$, the potential is finite and given by $e^{2}/4\pi a$ at lowest-order approximation, which is a feature of GQED. For larger distances $r\gg a$, we obtain the Coulomb potential as in QED in (3+1)D. On the other hand, for intermediate values $r\approx a$, we have a Coulomb potential with a screened electric charge. In Fig. 1, we plot Eq. (4) using an artificial value for $a$. Units are arbitrary for illustration. These results mostly concern the classical physics of GQED. From now on, we shall consider quantum effects of the dimensional reduction of GQED.
@@ II.1 The Feynman Rules
In this section, we derive the Feynman rules for PGQED in Eq. (II) at tree level. This is mandatory to apply the so-called Schwinger-Dyson equations for calculating the full Green functions of the model, which we shall discuss later. After a standard algebra, we obtain the bare gauge-field propagator, given by
@@
where we use the Landau gauge, i.e, $\xi\to+\infty$. After integrating out the gauge field in Eq. (II), it is straightforward to conclude that static charges interact through the potential $V(r)$ in Eq. (4), which is obtained from the Fourier transform of $\Delta^{0}_{00}(p_{0}=0,\textbf{p})$ [23]. Note that whether $\mu\rightarrow\infty$ (and equivalently $a\rightarrow 0$), we recover the PQED model as expected. Furthermore, in contrast to standard GQED, we can not conclude that this propagator in Eq. (5) describes either a massive or massless particle. Indeed, due to the square root in the gauge-field propagator, it does not have a pole and it needs a branch cut in order to become a single-valued function in the complex plane of $p$. Some important features of this propagator have been discussed in Ref. [20] at classical level.
@@ IV RAINBOW-QUENCHED APPROXIMATION
which is a differential equation for the mass function $\Sigma(p)$. This must also satisfy both the IR and UV conditions, given by Eqs. (29) and (30), respectively. Even after our set of approximations, Eq. (31) still is a nonlinear differential equation due to the denominator $p^{2}+\Sigma^{2}(p)$. This is a typical property of the Schwinger-Dyson equations, where the full propagator is written as a function of itself. Next, we propose a few approximations in order to circumvent this problem.
@@ IV.1 Approximations: $p\to 0$
$\frac{\pi}{2\alpha}$ $=\frac{1}{2}\ln{\left(\frac{M^{2}+\Lambda^{2}}{M^{2}}\right)}-\ln{\left(\frac{\sqrt{\mu^{2}+\Lambda^{2}}+\Lambda}{\mu}\right)}+$
$+\frac{|M|}{\sqrt{\mu^{2}-M^{2}}}\arctan{\left(\frac{\Lambda\sqrt{\mu^{2}-M^{2}}}{|M|\sqrt{\mu^{2}+\Lambda^{2}}}\right)}.$ (33)
@@
$\frac{\pi}{2\alpha}\approx\ln{\frac{\Lambda}{|M|}}-\ln{\left(\frac{\sqrt{\mu^{2}+\Lambda^{2}}+\Lambda}{\mu}\right)},$ (34)
@@ VI SUMMARY AND DISCUSSIONS
Recently, Ref. [20] performed a dimensional reduction of Lee-Wick electrodynamics from (3+1)D to (2+1)D. This calculation yields the same reduced gauge-field propagator we have obtained here, given by Eq. (5) in momentum space, although the starting action in (3+1)D is slightly different from ours. However, they are dual because the electrons interact through the same static potential $V(r)\propto(1-e^{-\mu r})/r$. In Ref. [20], it is also shown that this kind of model respects the microcausality at tree level, which is a desirable feature for a well-defined field theory. Here, our main results depend on the strength of quantum correction and are likely to be the same in the context of the reduced Lee-Wick electrodynamics, given the similarity of its gauge-field propagator in comparison to the one in PGQED. Obviously, our main results obtained here only hold for an initially massless Dirac field.
@@ Acknowledgement
The authors thank D. T. Alves, J. A. Helayel, J. D. L. Silva, R. Thibes, and M. Neves for interesting discussions about our work. C. A. P. C. J. is partially supported by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior Brasil (CAPES), finance code 001. L. O. N. and V. S. A. are partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico-Brazil (CNPq-Brazil), process: 408735/2023-6 CNPq/MCTI.